# Illuminating Black Hole Shadow with Dark Matter Annihilation

### v2

@@ I.5 Numerical Procedure and Results
For each point $\vec{r}$, we calculate the $\chi$ and $\beta$ functions along the streamline passing through it until the streamline reaches the boundary $s_{0}$. We consider the regions for the jet and the disk as follows:
[…]
Due to the existence of the stagnation surface, the upper limit in the jet region is smaller than that in the disk. The range for $\theta$ is determined by the asymptotic behavior that characterizes the boundary between the jet and disk regions , with $\theta$ approaching $\pi/2$ near the BH event horizon and $\pi/4$ at spatial infinity.
Figure S2: Examples of the electron-positron energy spectra, ${\rm d}n_{e}/{\rm d}\log_{10}p\equiv 4\pi\,(\ln 10)\,p^{3}\,f_{e}(\vec{r},p)$ for both the $b\bar{b}$ and $e^{+}e^{-}$ channels, at various benchmark DM masses. The spectra are depicted with solid lines for the location in the jet region $(r,\theta)=(4\,r_{g},\pi/16)$ and with dashed lines for the point in the disk region $(r,\theta)=(4\,r_{g},7\pi/16)$.
@@
Incorporating Eq. (S35) and Eq. (S12) into Eq. (S9), we obtain $f_{e}(\vec{r},p)$. The range for $p_{\rm inj}$ is specified by the injection spectrum ${\rm d}N^{\rm inj}_{e^{\pm},i}/{\rm d}E\,(p_{\rm inj})$, generated using the PPPC Cirelli et al. (2011) and MadDM Ambrogi et al. (2019) packages, with the latter applied exclusively to the $e^{+}e^{-}$ channel for DM masses below 55 55 GeV. Furthermore, the injection spectrum for the $e^{+}e^{-}$ channel exhibits a sharp peak as $p_{\rm inj}$ approaches $m_{\rm DM}$. To improve numerical precision during integration, we fit the region near the peak with a half-Gaussian function, ensuring normalization to maintain the total number of injected $e^{\pm}$ particles. Based on the given injection spectra, the range of $p_{\rm inj}$ is determined as follows:
$\displaystyle p_{\rm inj}\in\begin{cases}[1\,{\rm keV},\,m_{\rm DM}]&\quad\text{for}\,\,m_{\rm DM}\in[1\,{\rm MeV},\,5\,{\rm GeV})\,,\\ [10^{-7}\,m_{\rm DM},\,m_{\rm DM}]&\quad\text{for}\,\,m_{\rm DM}\in[5\,{\rm GeV},\,6\,{\rm TeV}]\,.\end{cases}$ (S39)
[…]
@@
In Fig. S2, we present the electron-positron energy spectra ${\rm d}n_{e}/{\rm d}\log_{10}p\equiv 4\pi\,(\ln 10)\,p^{3}\,f_{e}(\vec{r},p)$ for both the $e^{+}e^{-}$ and $b\bar{b}$ channels, across various benchmark DM masses. We have selected spatial positions $\vec{r}$ corresponding to two distinct points: one in the jet region $(r,\theta)=(4\,r_{g},\pi/16)$ contributing to emissions in the inner shadow, and another in the disk region $(r,\theta)=(4\,r_{g},7\pi/16)$. The distribution of the electron-positron energy spectrum generally maintains the shape of the injection spectrum, although distortions and energy shifts occur due to propagation effects. Specifically, in the jet region, the adiabatic compression rate $\dot{p}_{\rm adi}$ can surpass the synchrotron loss rate $\dot{p}_{\rm syn}$, consequently accelerating electrons and positrons to energies exceeding $m_{\rm DM}$. Conversely, in the disk region, adiabatic expansion dominates, leading to energy losses that predominantly shape the propagation process and truncate the energy spectrum around $p\sim m_{\rm DM}$.
Figure S3: The electron-positron number density, $n_{e}(\vec{r}\,)\equiv\textstyle\int 4\pi p^{2}f_{e}(\vec{r},p)\,{\rm d}p$, depicted across the $(m_{\rm DM},\,\,\langle\sigma v\rangle)$ plane for both the $b\bar{b}$ and $e^{+}e^{-}$ channels. The spatial locations are consistent with those in Fig. S2.
@@
  • When $\langle\sigma v\rangle\ll\langle\sigma v\rangle_{\rm sat}$, the nearby DM energy density conforms to the spike profile $\rho(\vec{r}\,)=\rho_{0}\,(r/r_{0})^{-\gamma_{\rm sp}}$. Therefore, $n_{e}(\vec{r})$ varies as $\langle\sigma v\rangle/m_{\rm DM}^{2}$, aligning with the contour slope in the low $\langle\sigma v\rangle$ region depicted in Fig. S3.
@@ II Covariant Radiative Transfer and Intensity Map
Following the computation of the electron and positron spectrum, $f_{e}(\vec{r},p)$, we employ the covariant radiative transfer formalism Gammie and Leung (2012); Dexter (2016) to determine the intensity at each point on the observer plane, contrasting these results with the astrophysical background using the RAPTOR package Bronzwaer et al. (2018, 2020). This section elaborates on the methodology.
To compute the flux, we initially calculate geodesics that connect the observer to the BH, utilizing the backward ray tracing method Bardeen et al. (1972); Luminet (1979). Each point on the observer plane is associated with a distinct initial direction of these geodesics. The observer plane’s critical curve demarcates the regions for geodesics that either propagate to infinity or terminate at the BH Podurets (1965); Ames and Thorne (1968); Luminet (1979).
The next step involves integrating emissions along the geodesics, tracing from the BH side towards the observer. A practical approach for this integration is through the covariant radiative transfer equation Gammie and Leung (2012); Dexter (2016):
@@
defined within a local reference frame. In this equation, $\lambda$ denotes the affine parameter, $I_{\nu}$ the intensity, $j_{\nu}$ the emissivity, and $\alpha_{\nu}$ the absorption coefficient, with the subscript $\nu$ indicating the photon frequency in the reference frame, which is related to the observed frequency $\nu_{0}$ by a redshift factor. The calculation can proceed in any frame as the quantities $I_{\nu}/\nu^{3}$, $j_{\nu}/\nu^{2}$, and $\nu\alpha_{\nu}$ are invariant under Lorentz transformation.
@@
[…]
@@
aligning with observations that favor an optically thin accretion flow at $\nu_{0}=230$ GHz Yuan and Narayan (2014); Prieto et al. (2016); et al (2019).
@@
where $\beta_{p}\equiv P_{\text{gas}}/P_{\text{mag}}$ represents the ratio of gas pressure to magnetic pressure, with $P_{\text{mag}}$ being defined as $B^{2}/2$. Here, $R_{\text{high}}$ and $R_{\text{low}}$ represent the temperature ratios of protons to electrons in areas of high and low magnetic field strengths, respectively. $R_{\text{low}}$ is set to 11 11 in the simulation, in accordance with Refs. et al (2019); Akiyama et al. (2021b, 2023). Variations in $R_{\text{low}}$ have been found to exert quantitatively negligible effects on the intensity distribution Mizuno et al. (2021); Zhang et al. (2024).
In the maintext, we present results specifically for a high reflectivity value of $R_{\text{high}}=1$ and a spin parameter $a_{J}=0.9375$, consistent with observations by the EHT et al (2019); Akiyama et al. (2021b, 2023). To ensure the robustness of our constraints, we also explore a range of these parameters, varying $R_{\text{high}}$ from 11 11 to 160160 160160 and adjusting $a_{J}$ down to 0.50.5 0.50.5 Cruz-Osorio et al. (2022). For each scenario, we maintain the total intensity of the BH image at 0.60.6 0.60.6 Jy, which serves to determine the normalization of the GRMHD profile alongside the electron temperature $T_{e}$. Our analysis reveals that these variations have a quantitatively minor impact, altering the intensity by less than an order of magnitude.
The exclusion region for the $e^{+}e^{-}$ annihilation channel displays a notable shift for DM masses below GeV scales, differing significantly from the higher mass region and the $b\bar{b}$ channel, as illustrated in Fig. 3 of the maintext. Specifically, the $b\bar{b}$ channel excludes regions of highest $n_{e}$ shown in Fig. S3, whereas the $e^{+}e^{-}$ channel exhibits a distinct turning point around 0.10.1 0.10.1 GeV. This shift is attributed to a pronounced drop in the synchrotron radiation flux at 230230 230230 GHz for DM masses below 0.10.1 0.10.1 GeV. Accounting for gravitational redshift near an emission point approximately $4\,r_{g}$ from the SMBH, the photon frequency in the plasma frame is about 300300 300300 GHz. The synchrotron radiation spectrum, for an electron or positron with momentum $p$, adheres to the $F$ function as defined in Eq. (S42). This function presents a gentle slope $F\sim(\nu/\nu_{c})^{1/3}$ for $\nu/\nu_{c}\ll 1$, transitioning to a steep decline $F\sim\exp(-\nu/\nu_{c})$ for $\nu$ exceeding $\nu_{c}$. The critical frequency $\nu_{c}\equiv 3eBp^{2}\sin{\theta_{B}}/(4\pi m_{e}^{3})$ is estimated numerically as:
@@
Referencing Fig. S2, the spectrum of electrons and positrons predominantly corresponds to $p\sim m_{\rm DM}$. Thus, most synchrotron radiation emanating from the annihilation products of DM with mass below 0.10.1 0.10.1 GeV fails to contribute significantly at $\nu=230$ GHz.
@@ III Critical Curve and Inner Shadow Contour
In this section, we elucidate the equations defining the critical curve and the contour of the inner shadow, as depicted in Fig. 2, employing dimensionless notation denoted by $\tilde{}\,$, normalized by the gravitational radius $r_{g}$.
We commence by defining a universal coordinate system $(\tilde{\alpha},\tilde{\beta})$ on the observer plane, where the $\tilde{\alpha}$-axis is perpendicular to, and the $\tilde{\beta}$-axis is parallel with, the projected spin of the BH Bardeen et al. (1972); Luminet (1979). Each point on this plane corresponds to a null geodesic in Kerr spacetime, described by two conserved quantities: the energy-rescaled angular momentum $\tilde{\lambda}$ and the Carter constant $\tilde{\eta}$, expressed as Carter (1968); Gralla and Lupsasca (2020a, b); Johnson et al. (2020); Lupsasca et al. (2024):
$\begin{split}\tilde{\lambda}&=-\tilde{\alpha}\sin\theta_{o},\\ \tilde{\eta}&=(\tilde{\alpha}^{2}-a_{J}^{2})\cos^{2}\theta_{o}+\tilde{\beta}^{2},\end{split}$ (S49)
@@
The critical curve demarcates the boundary on the observer plane between geodesics that terminate on the BH (inner) and those that escape to infinity (outer). It signifies the locus of bound orbits around the Kerr BH Podurets (1965); Ames and Thorne (1968); Luminet (1979); Falcke et al. (2000); Gralla et al. (2019), maintaining a constant radius $\tilde{r}_{c}$ within the range $\tilde{r}_{c}\in[\tilde{r}_{c}^{-},\tilde{r}_{c}^{+}]$, where
$\tilde{r}_{c}^{\pm}\equiv 2\left[1+\cos\left(\frac{2}{3}\arccos(\pm a_{J})\right)\right].$ (S50)
Determining the bound orbits involves identifying the double roots of the radial potential in the geodesics equation, leading to:
$\begin{split}\tilde{\lambda}_{c}&=a_{J}+\frac{\tilde{r}_{c}}{a_{J}}\left[\tilde{r}_{c}-\frac{2\Delta(\tilde{r}_{c})}{\tilde{r}_{c}-1}\right],\\ \tilde{\eta}_{c}&=\frac{\tilde{r}_{c}^{3}}{a_{J}^{2}}\left[\frac{4\Delta(\tilde{r}_{c})}{(\tilde{r}_{c}-1)^{2}}-\tilde{r}_{c}\right],\end{split}$ (S51)
where $\Delta(\tilde{r}_{c})=\tilde{r}_{c}^{2}+a_{J}^{2}-2\,\tilde{r}_{c}$. By parameterizing $\tilde{r}_{c}$ and correlating Eq. (S49) with Eq. (S51), we formulate a parametric depiction of the critical curve on the $\tilde{\alpha}-\tilde{\beta}$ plane.
@@
$\tilde{\rho}=\left[2\sqrt{\tilde{r}_{H}}+\left(1+\frac{1}{2}\cos^{2}\theta_{o}\right)\arctan\left[\sin\varphi\tan\theta_{o}\right]\right],$ (S52)

### v3

@@ I.5 Numerical Procedure and Results
Figure S2: Examples of the electron-positron energy spectra, ${\rm d}n_{e}/{\rm d}\log_{10}p\equiv 4\pi\,(\ln 10)\,p^{3}\,f_{e}(\vec{r},p)$ for both the $b\bar{b}$ and $e^{+}e^{-}$ channels, at various benchmark DM masses. Solid lines represent spectra at a typical jet location $(r,\theta)=(4\,r_{g},\pi/16)$, while dashed lines correspond to a representative disk location $(r,\theta)=(4\,r_{g},7\pi/16)$. In most cases, synchrotron losses dominate over adiabatic acceleration. The $b\bar{b}$ channel yields softer spectra compared to the $e^{+}e^{-}$ channel, reflecting differences in the injection spectra. For heavier DM, stronger synchrotron cooling leads to greater spectral suppression. For $m_{\rm DM}=0.1\,$GeV, the calculation terminates at non-relativistic momenta $p\sim\mathrm{MeV}$.
For each point $\vec{r}$, we compute the $\chi$ and $\beta$ functions along the corresponding streamline up to the boundary point $s_{0}$. Substituting Eqs. (S35) and (S12) into Eq. (S9) yields the phase space density $f_{e}(\vec{r},p)$. In both the jet and disk regions, our calculations are confined to the domain specified in Sec. I.3, with $r\in[2.5\,r_{g},30\,r_{g}]$. For locations within this radial range but lying outside the defined jet and disk regions, we interpolate $\log_{10}f_{e}(\vec{r},p)$ linearly along the $\theta$ direction at fixed $r$. In the inner region where $r<2.5\,r_{g}$, we set $f_{e}(\vec{r},p)$ equal to its value at $r=2.5\,r_{g}$ for the corresponding $\theta$.
@@
The range for $p_{\rm inj}$ is specified by the injection spectrum ${\rm d}N^{\rm inj}_{e^{\pm},i}/{\rm d}E\,(p_{\rm inj})$, generated using the PPPC Cirelli et al. (2011) and MadDM Ambrogi et al. (2019) packages, with the latter applied exclusively to the $e^{+}e^{-}$ channel for DM masses below $5$ GeV. Furthermore, the injection spectrum for the $e^{+}e^{-}$ channel exhibits a sharp peak as $p_{\rm inj}$ approaches $m_{\rm DM}$. To improve numerical precision during integration, we fit the region near the peak with a half-Gaussian function, ensuring normalization to maintain the total number of injected $e^{\pm}$ particles. Based on the given injection spectra, the range of $p_{\rm inj}$ is determined as follows:
$\displaystyle p_{\rm inj}\in\begin{cases}[1\,{\rm keV},\,m_{\rm DM}]&\quad\text{for}\,\,m_{\rm DM}\in[1\,{\rm MeV},\,5\,{\rm GeV})\,,\\ [10^{-8}\,m_{\rm DM},\,m_{\rm DM}]&\quad\text{for}\,\,m_{\rm DM}\in[5\,{\rm GeV},\,10\,{\rm TeV}]\,.\end{cases}$ (S38)
Consequently, the range of $p$ is set as $p\in[1\,{\rm MeV},\,10\,{\rm TeV}]$, where the lower bound of $p$ is truncated at $1$ MeV to ensure that electrons and positrons remain relativistic. This broader range of $p$ compared to $p_{\rm inj}$ allows for the inclusion of both acceleration and deceleration processes that electrons and positrons undergo during propagation.
@@
In Fig. S2, we present the electron-positron energy spectra ${\rm d}n_{e}/{\rm d}\log_{10}p\equiv 4\pi\,(\ln 10)\,p^{3}\,f_{e}(\vec{r},p)$ for both the $e^{+}e^{-}$ and $b\bar{b}$ channels, across various benchmark DM masses. We have selected spatial positions $\vec{r}$ corresponding to two distinct points: one in the jet region $(r,\theta)=(4\,r_{g},\pi/16)$ contributing to emissions in the inner shadow, and another in the disk region $(r,\theta)=(4\,r_{g},7\pi/16)$. The distribution of the electron-positron energy spectrum generally maintains the shape of the injection spectrum, although distortions and energy shifts occur due to propagation effects.
Figure S3: Electron-positron number density, $n_{e}(\vec{r}\,)\equiv\textstyle\int 4\pi p^{2}f_{e}(\vec{r},p)\,{\rm d}p$, produced from DM annihilation after propagation, shown across the $(m_{\rm DM},\,\langle\sigma v\rangle)$ parameter space for both the $b\bar{b}$ and $e^{+}e^{-}$ channels. The density is evaluated at representative points in the jet region $(r,\theta)=(4\,r_{g},\pi/16)$ and the disk region $(r,\theta)=(4\,r_{g},7\pi/16)$, consistent with Fig. S2. The horizontal contours arise from the formation of a central core in the DM profile at large cross sections, while the tilted contours reflect the steeper spike regime at lower cross sections.
@@
  • When $\langle\sigma v\rangle\ll\langle\sigma v\rangle_{\rm sat}$, the nearby DM energy density conforms to the spike profile $\rho(\vec{r}\,)=\rho_{0}\,(r/r_{0})^{-7/3}$. Therefore, $n_{e}(\vec{r}\,)$ varies as $\langle\sigma v\rangle/m_{\rm DM}^{2}$, aligning with the contour slope in the low $\langle\sigma v\rangle$ region depicted in Fig. S3.
@@ II Covariant Radiative Transfer and Intensity Map
Following the computation of the electron-positron spectrum, $f_{e}(\vec{r},p)$, we employ the covariant radiative transfer formalism Gammie and Leung (2012); Dexter (2016) to determine the intensity at each point on the observer plane and compare it with the astrophysical background. This section details the methodology, utilizing the RAPTOR package Bronzwaer et al. (2018, 2020) for numerical calculations.
To compute the flux, we first solve for geodesics that connect the observer to the BH using the backward ray tracing method Bardeen et al. (1972); Luminet (1979). Each point on the observer plane corresponds to a unique initial geodesic direction. The critical curve on the observer plane delineates regions where geodesics either escape to infinity or terminate at the BH Podurets (1965); Ames and Thorne (1968); Luminet (1979).
Next, we integrate emissions along these geodesics, tracing from the BH toward the observer. This integration follows the covariant radiative transfer equation Gammie and Leung (2012); Dexter (2016):
@@
which is formulated in a local reference frame. Here, $\lambda$ is the affine parameter along the geodesics, $I_{\nu}$ is the intensity, $j_{\nu}$ is the emissivity, and $\alpha_{\nu}$ is the absorption coefficient. The subscript $\nu$ denotes the photon frequency in the local frame, which relates to the observed frequency $\nu_{0}$ via a redshift factor. The computation can be performed in any frame since the quantities $I_{\nu}/\nu^{3}$, $j_{\nu}/\nu^{2}$, and $\nu\alpha_{\nu}$ remain invariant under Lorentz transformations.
@@
$j_{\nu}=\int f_{e}(\vec{r},p)\,\eta\,4\pi p^{2}\,{\rm d}p,\qquad\eta=\frac{\sqrt{3}e^{3}B\sin\theta_{B}}{8\pi^{2}m_{e}}F\left(\frac{\nu}{\nu_{c}}\right),\qquad F(x)\equiv x\int_{x}^{\infty}K_{5/3}(\zeta)\,{\rm d}\zeta,$ (S41)
@@
aligning with observations that favor an optically thin accretion flow at $\nu_{0}=230$ GHz Yuan and Narayan (2014); Prieto et al. (2016); Akiyama et al. (2019c).
@@
where $\beta_{p}\equiv P_{\text{gas}}/P_{\text{mag}}$ represents the ratio of gas pressure to magnetic pressure, with $P_{\text{mag}}$ being defined as $B^{2}/2$. Here, $R_{\text{high}}$ and $R_{\text{low}}$ represent the temperature ratios of protons to electrons in areas of high and low magnetic field strengths, respectively. In the maintext, we adopt fiducial values of $R_{\text{high}}=R_{\text{low}}=1$, consistent with those supported by EHT observations Akiyama et al. (2019c, 2021b, 2023). In the final section, we examine potential uncertainties in these parameters based on both observational data and two-temperature GRMHD simulations, and show that their impact on our constraints is minor.
The exclusion region for the $e^{+}e^{-}$ annihilation channel displays a notable shift for DM masses below GeV scales, differing significantly from the higher mass region and the $b\bar{b}$ channel, as illustrated in Fig. 3 of the maintext. Specifically, the $b\bar{b}$ channel excludes regions of highest $n_{e}$ shown in Fig. S3, whereas the $e^{+}e^{-}$ channel exhibits a distinct turning point around $0.1$ GeV. This shift is attributed to a pronounced drop in the synchrotron radiation flux at $230$ GHz for DM masses below $0.1$ GeV. Accounting for gravitational redshift near an emission point approximately $4\,r_{g}$ from the SMBH, the photon frequency in the plasma frame is about $300$ GHz. The synchrotron radiation spectrum, for an electron or positron with momentum $p$, adheres to the $F$ function as defined in Eq. (S41). This function presents a gentle slope $F\sim(\nu/\nu_{c})^{1/3}$ for $\nu/\nu_{c}\ll 1$, transitioning to a steep decline $F\sim\exp(-\nu/\nu_{c})$ for $\nu$ exceeding $\nu_{c}$. The critical frequency $\nu_{c}\equiv 3eBp^{2}\sin{\theta_{B}}/(4\pi m_{e}^{3})$ is estimated numerically as:
@@
Referencing Fig. S2, the spectrum of electrons and positrons predominantly corresponds to $p\sim m_{\rm DM}$. Thus, most synchrotron radiation emanating from the annihilation products of DM with mass below $0.1$ GeV fails to contribute significantly at $\nu=230$ GHz.
@@ III Critical Curve and Inner Shadow Contour
In this section, we elucidate the equations defining the critical curve and the contour of the inner shadow, as depicted in Fig. 2.
We start by defining a universal coordinate system $({\alpha},{\beta})$ on the observer plane, where the ${\alpha}$-axis is perpendicular to, and the ${\beta}$-axis is parallel with, the projected spin of the BH Bardeen et al. (1972); Luminet (1979). The impact parameters $\alpha$ and $\beta$ are dimensionless, normalized by the gravitational radius $r_{g}$.
Each point on this plane corresponds to a null geodesic in Kerr spacetime, described by two conserved quantities: the energy-rescaled angular momentum ${\lambda}$ and the Carter constant ${\eta}$, expressed as Carter (1968); Gralla and Lupsasca (2020a, b); Johnson et al. (2020); Lupsasca et al. (2024):
$\begin{split}{\lambda}&=-{\alpha}\sin\theta_{o},\\ {\eta}&=({\alpha}^{2}-a_{J}^{2})\cos^{2}\theta_{o}+{\beta}^{2},\end{split}$ (S48)
@@
The critical curve delineates the boundary on the observer plane between geodesics that terminate on the BH (inner region) and those that escape to infinity (outer region). The corresponding geodesics form bound orbits around the Kerr BH Podurets (1965); Ames and Thorne (1968); Luminet (1979); Falcke et al. (2000); Gralla et al. (2019), maintaining a constant radius $\mathcal{R}_{c}$ within the range $\mathcal{R}_{c}\in[\mathcal{R}_{c}^{-},\mathcal{R}_{c}^{+}]$, where
$\mathcal{R}_{c}^{\pm}\equiv 2\left[1+\cos\left(\frac{2}{3}\arccos(\pm a_{J})\right)\right].$ (S49)
The determination of bound orbits requires identifying the double roots of the radial potential in the geodesic equation, leading to
$\begin{split}{\lambda}_{c}&=a_{J}+\frac{\mathcal{R}_{c}}{a_{J}}\left[\mathcal{R}_{c}-\frac{2\Delta(\mathcal{R}_{c})}{\mathcal{R}_{c}-1}\right],\\ {\eta}_{c}&=\frac{\mathcal{R}_{c}^{3}}{a_{J}^{2}}\left[\frac{4\Delta(\mathcal{R}_{c})}{(\mathcal{R}_{c}-1)^{2}}-\mathcal{R}_{c}\right],\end{split}$ (S50)
where $\Delta(\mathcal{R}_{c})=\mathcal{R}_{c}^{2}+a_{J}^{2}-2\,\mathcal{R}_{c}$. By equating Eq. (S48) with Eq. (S50), we construct a parametric representation of the critical curve in the ${\alpha}-{\beta}$ plane for different values of $\mathcal{R}_{c}$.
@@
${\rho}=\left[2\sqrt{\mathcal{R}_{H}}+\left(1+\frac{1}{2}\cos^{2}\theta_{o}\right)\arctan\left[\sin\varphi\tan\theta_{o}\right]\right],$ (S51)